Write the equation for the line that is (a)parallel and (a) perpendicular to y= 8\7x+6 passing through (-2,2)
Answer by mananth(16946)

PARALLEL
y = 8/7 x + 6
Compare this equation with y=mx+b
slope m = 8/7 0.8
The slope of a line parallel to the above line will be the same
The slope of the required line will be 8/7
m= 8/7 ,point ( -2 , 2 )
Find b by plugging the values of m & the point in
y=mx+b
2 = -16/7 + b
b= 30/7

m= 8/7
Plug value of the slope and b in y = mx +b
The required equation isy = 8/ 7 x + 30/ 7
PERPENDICULAR
1 y = 8/7 x + 6


Compare this equation with y=mx+b
slope m = 8/7

The slope of a line perpendicular to the above line will be the negative reciprocal -7/8
m1*m2=-1
The slope of the required line will be -7/8

m= - 7/8 ,point ( -2 , 2 )
Find b by plugging the values of m & the point in
y=mx+b
2 = 7/4 + b
b= 1/4
m= - 7/8
The required equation is y = - 7/ 8 x + 1/ 4


m.ananth@hotmail.ca